Suppose you have $40,000 to invest. Youre considering Miller-Moore Equine Enterprises (MMEE), which is currently selling for $40 per share. You notice that a put option with a $40 strike is available with a premium of $3.20. Calculate your percentage return on the put option for the six-month holding perlod if the stock price declines to $36 per share. (A negative value should be indicated by a minus sign. Leave no cells blank-be certain to enter O wherever required. Do not round intermediate calculations. Enter your 6 month return as a percent rounded to 2 decimal places.) Percentage return
Answer #1

Answer = 25.00%

Put premium = 3.2
Payoff from put = (40-36) 4
Gain from put = (4-3.2) 0.8
return = 0.8/3.2 25.00%